OpenStudy (anonymous):

4(3-x)4/3-5=99 A. –5, 11 B. 5 C. 11 D. –11

OpenStudy (mayankdevnani):

@tonirenee_ i think this question need the value of x. right???

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

4 (3-x)^4/3-5=99

OpenStudy (mayankdevnani):

\[\huge 4(3-x) \times \frac{4}{3} -5=99\] Is this your equation?????

OpenStudy (anonymous):

no its 4(3-x)^4/3-5=99

OpenStudy (anonymous):

\[4 (3-x)^{4/3}-5=99\]

OpenStudy (mayankdevnani):

oh!! i see

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

someone help plz